OpenStudy (anonymous):

i get how 27 mod is 12 = 1, but how is -27 mod 12 = 9. show your steps please :)

OpenStudy (anonymous):

i mean 27 mod 12 = 1

OpenStudy (anonymous):

should be 3--- 27mod12 = 3

OpenStudy (anonymous):

3 is the remainder from the division of 27/12

OpenStudy (anonymous):

works like a clock ... only this time you're going in the opposite direction look at a watch

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=-27+mod+12

OpenStudy (anonymous):

means remainder when you divide -27 by 12 is 9, or \[-27=9 - 3\times 12\]

OpenStudy (anonymous):

\[x\equiv y ( \text{ mod }n)\] means \[n|(x-9)\] and so \[12|-27-9=-36\]

OpenStudy (anonymous):

thanks satellite!!!!! i get it now! :)